Where applicable, and if not specifically addressed in the question, assume interest is compounded annually ( i.e. P/Y = 1) , at the end of each year.  The only exception will be for the mortgage problem (#34), which will have payments and interest calculated on a monthly basis, at the end of each month.
Answer #1

A is the correct answer

Monthly payment can be calculated using PMT function in excel or calculator

N = 30 x 12, PV = 300,000, FV = 0, I/Y = 6%/12

=> Compute PMT = $1,798.65 is monthly payment

Total Principal Paid over 12 years can be calculated using CUMPRINC function in excel

Principal Paid = CUMPRINC(rate = 6%/12, nper = 30 x 12, pv = 300,000, 1, 12 x 12, 0) = $62,761.68

Principal Outstanding = 300,000 - 62,761.68 = 237,238.32

Total Interest Paid = Total Payment - Principal Paid

= 1,798.65 x 12 x 12 - 62,761.68

= 196,244.15